Compute (a) return on equity (ROE), (b) market-to- book (M/B) ratio, and (c) enterprise value based on the following information: net profit margin = 4%, asset turnover = 1.5, debt = $60 million (assuming that the book value and market value of debt are the same), book value of assets = $100 million, market capitalization = $80 million, cash = $20 million.
Answer #1

Hi

Profit Margin = 4%

asset turnover = 1.5 times

Debt = 60 million

Cash = $20 million

Book Value of Asset = $100 million

Market Cap = $80 million

So Total Sales = asset turnover* Asset = 1.5*100 = $150 million

Net Income = 4%*150 = $6 million

Equity = Asset - Debt = 100 - 60 = $40 miilion

ROE = Net Income/ Equity

= 6/40

= 15%

Market to Book Ratio = Market Cap/ Book Value of Equity

= 80/40 = 2.0 times

Enterprise Value EV = Debt + Market Cap - Cash

= 60 +80 - 20

= $120 million

Thanks